# Defining a Role-Centered Terminology for Physical Representations and Controls

### v1

@@ 2.2. Blended Words
Linguistics studies the evolution of everyday language over time, including word-formation processes (e.g., derivations, abbreviations, and acronyms). Some new words are created by combining splinters (Barrena Jurado, 2019) excerpted from source words. For example, blended words concatenate a basis from a first word and a suffix from a second word, with phonemic overlap (Beliaeva, 2016; Connolly, 2013; Gries, 2004). Linguists observed several possibilities and strategies of splinters excerption and combination that enter the definition of “blended words” that keep source words recognisable (Gries, 2004, 2006). For instance, “heliport” (heli(copter) + (air)port), “motel” (mot(or) + (h)otel), and “brunch” (br(eakfast) + (l)unch) are blend words (Connolly, 2013; Gries, 2004). However, blends can follow some other word-formation processes. For example, clipping compounds concatenate the bases of two words (Beliaeva, 2016; Gries, 2004, 2006) (e.g., “fintech” (fin(ancial) tech(nology)), “sysadmin” (sys(tem) + admin(istrator)), and “hydrail” (hyd(rogen) + rail(way))). Word blending enables creating new words of only 1, 2, or 3+ syllables (Connolly, 2013), thus creating new vocabulary that designates concepts through single words, which also facilitates communication.
Blended words already exist in Computer Science and Human–Computer Interaction: for example, “pixel” (pic(ture)s + el(ement)) and “widget” (wi(ndow) + (ga)dget) are kinds of blended words that have become common in the usual language and are referenced in standard dictionaries. In the field of tangible user interfaces, several of the aforementioned terms are blended words, such as “phicon” (ph(ysical) + icon) (Ullmer and Ishii, 1997), “phandle” (ph(ysical) + (h)andle) (Ullmer and Ishii, 1997), and “phidget” (ph(ysical) + (w)idget) (Greenberg and Fitchett, 2001).
@@ 2.3. Taxonomies and Classifications
The literature has resorted to taxonomies and classifications to characterize the conceptual space of physical user interfaces. Whereas the seminal work on graspable user interfaces proposed a conceptual space with thirteen axes (Fitzmaurice et al., 1995), Ishii later classified tangible user interfaces through eight genres (Ishii, 2008). About two dozen frameworks organized the conceptual space along different numbers of axes Mazalek and Hoven, 2009, Shaer and Hornecker, 2010: Ch. 5, Ullmer et al., 2022: Ch. 3.
For example, Fishkin’s taxonomy of physicality (or tangibility) (Fishkin, 2004) organized physical user interfaces as a twenty-item conceptual space along two axes: embodiment (i.e., the physical relationship between the input and the output of interaction loops) and metaphor (i.e., the digital meaning given to interaction loops with physical representations).
All those frameworks were organized by types: abstracting, designing, and building; and by facets: experiences, domains, physicality, interactions, and technologies (Mazalek and Hoven, 2009).
@@ 3.2. Back to Tangible User Interfaces’ Concept
According to Ullmer and Ishii: “Tangible interfaces give physical form to digital information, employing physical artifacts both as representations and controls for computational media” (Ullmer and Ishii, 2000). This definition’s phrasing reveals Ullmer and Ishii’s datum-centered viewpoint on tangible user interfaces (Shaer and Hornecker, 2010: p. 19). The adjective “tangible” should then be used to qualify data that becomes tangible, not to qualify artifacts or bodied objects (the latter is already “tangible” in the usual sense). This article claims we should then rather talk about “tangible data,” and no more about “tangible artifacts” or “tangible objects.” Thus, tangible data are data that become tangible through artifacts (that give physical form to data): only those “pure” artifacts must be qualified as tangible representations to match Ullmer and Ishii’s definition strictly. By extending this line of reasoning, this article proposes naming all artifacts according to their digital role—as instantiations of digital information—by providing meaning about how they represent digital entities. This way, the following section defines a what–how terminology to name physical representations and controls from a role-centered viewpoint.
@@ 3.3. A Role-Centered Terminology
The new terminology uses blended words (Beliaeva, 2016; Connolly, 2013) formed from a taxonomy on “what” digital entities come into the digital world and “how”. The first part concerns what digital entities are represented or controlled by physical representations:
@@
    Operations involve triggering an event or action in the computing system (e.g., core operations (Ullmer et al., 2008) include loading or saving data, loading an app, and validation).
@@
    Finally, constraints structure the behavior of data representations, thus structuring mapped digital information. They can give meaning to artifacts by triggering actions or operations (e.g., detecting a block’s presence to trigger an action (Ullmer and Ishii, 1999; Ullmer et al., 1998)). Broadly, “constraints” affect physical objects’ positioning (e.g., table surface) before being sensed by the computer. “Constraints” are not always computer-generated: for example, some circles where to place artifacts can be video-projected onto a surface or drawn with paint.
@@
    Those that are made “graspable” are represented by symbolic artifacts of a generic form, with a low level of embodiment. Saying tangible or graspable is not saying better or worse, but simply different gradations of embodiment (Fishkin, 2004) that correspond to different design and technology answers to different needs.
@@
From the seven source words mentioned above, the seven following splinters (Barrena Jurado, 2019) are excerpted: four bases from what words: “dat-,” “tol-,” “op-,” and “const-”; and three suffixes from how words: “-ible,” “-able,” and “-nible.” Several attempts were required to select those splinters by a compromise between pronunciation, syllable count, and recognition of original words in the final terms (at least for experts aware of the terminology’s construction). Blending those bases and suffixes thus provides twelve terms embedding meaning about what and how digital entities come to the physical world (see Table 1):
  1. (1)
    Datible [”deItIb9l];
  2. (2)
    Datable [”deIt9b9l];
  3. (3)
    Datnible [”deIt.nIb9l];
  4. (4)
    Tolible [tu:lIb9l];
  5. (5)
    Tolable [tu:l9b9l];
  6. (6)
    Tolnible [tu:l.nIb9l];
  7. (7)
    Opible [OpIb9l];
  8. (8)
    Opable [Op9b9l];
  9. (9)
    Opnible [Op.nIb9l];
  10. (10)
    Constible [”kOnstIb9l];
  11. (11)
    Constable [”kOnst9b9l];
  12. (12)
    Constnible [”kOnst.nIb9l].
@@
Datum DATIBLE‘Datum is tangible’ DATABLE‘Datum is graspable’ DATNIBLE‘Datum is intangible’
Tool TOLIBLE‘Tool is tangible’ TOLABLE‘Tool is graspable’ TOLNIBLE‘Tool is intangible’
Operation OPIBLE‘Operation is tangible’ OPABLE‘Operation is graspable’ OPNIBLE‘Operation is intangible’
Constraint CONSTIBLE‘Constraint is tangible’ CONSTABLE‘Constraint is graspable’ CONSTNIBLE‘Constraint is intangible’
@@ 3.5. Four Role-Centered Tangibility Classes
  • In Class IV specimens, only operations are provided, through tangible or graspable entities (i.e., presence of opibles or opables, and data and tools are unknown or absent). This class provides specimens that are not restricted to a specific application and can thus control operations shared across applications (i.e., core operations (Ullmer et al., 2008)).
The application hallmarks’ patterns that correspond to these tangibility classes are reported in Table 2.
@@
The following section illustrates the descriptive power of the twelve terms through examples from the literature.
@@ 4.1. Specimens’ Selection
Because researchers have produced hundreds of tangible user interface specimens over the past few decades (Fleck et al., 2018), the present collection is not meant to be exhaustive, but a rather representative sample. The specimen selection process is neither as rigorous as that in systematic literature reviews nor intended to provide reliable statistics. Instead, the objective is to illustrate the terminology coverage and usefulness by describing a variety of emblematic tangible user interface specimens and related physical user interfaces. Because research interests vary over the years, restricting the analysis to a specific date range could have led to the exclusion of some tangible user interface kinds; therefore, the entire research period, from the premise to the more recent work, was taken into account. The selection thus gathered various genres [^3] Seven genres of tangible user interface applications were described in 2008 (Ishii, 2008): (1) tangible telepresence, (2) tangibles with kinetic memory, (3) constructive assembly, (4) tokens and constraints, (5) interactive surfaces—tabletop TUI, (6) continuous plastic TUI, (7) augmented everyday objects, and (8) ambient media, to which actuated TUI and shape-changing TUI must now be added. of tangible user interfaces that were studied during this entire period. Finally, the selection was filtered to retain work that showed real applications. Work merely demonstrating a new technology or a new kind of tangible user interface was not retained, as its naming could vary across applications. Indeed, knowing the mapping to the digital world is necessary because being physical is not enough to be a tangible representation: the same representation can be tangible in some applications but not in others.
The selection process is summarized in Figure 1. Several sources were browsed. First, a previous state of the art (Shaer and Hornecker, 2010) and some specimens already known to the author (e.g., some specimens that were experienced during demonstration sessions at conferences over time). Second, the projects page [^4] MIT Tangible Media Group’s projects page: https://tangible.media.mit.edu/projects/, last accessed 2022-08-03. of the MIT Tangible Media Group, which is certainly the most prolific laboratory in this field (accounts for 45.5% of the final collection’s provenance). Third, to obtain even more recent work, sessions with ‘tangible’ in their title from the last ten CHI conference proceedings (from 2013 to 2022 venues). Finally, the last ten TEI conference proceedings from 2013 to 2022 were also browsed. Specimens regarding remote computer-supported conditions were rejected.
@@
State of the art (Shaer and Hornecker, 2010)
@@
Figure 1. The selection process of applicative specimens of tangible user interfaces and related physical user interfaces.:
@@ 4.2. Corpus Description
1. Slot Machine 1976 (Perlman, 1976) and see Rows (procedure) - - - $\times$ - $\times$ - CONSTABLE
in (Ullmer et al., 2022: Ch. 1) Buttons (on each row) - - $\times$ - - $\times$ - OPABLE
@@ 1. Slot Machine 1976 (Perlman, 1976) and see Rows (procedure) - - - $\times$ - $\times$ - CONSTABLE
2. CAAD 3D Modelling System 1979 (Aish, 1979; Aish and Noakes, 1984) Building Blocks $\times$ - - - - $\times$ - DATABLE
@@
3. Self-Builder Model (Segal Model) 1980 (Frazer and Frazer, 1982, 1980) cited Panels $\times$ - - - $\times$ - - DATIBLE
from (Sutphen et al., 2000) Board - - - $\times$ - $\times$ - CONSTABLE
@@
4. Marble Answering Machine 1992 See in Follmer et al., 2013, Ishii and Ullmer,1997 Machine - - - $\times$ - $\times$ - CONSTABLE
and (Ullmer et al., 2022: Ch. 1) Machine’s indentation (play slot) - - $\times$ - - $\times$ - OPABLE
@@
5. Head Prop 1994 (Hinckley et al., 1994) Doll’s head (brain) $\times$ - - - $\times$ - - DATIBLE
@@
6. GraspDraw 1995 (Fitzmaurice et al., 1995) Bricks - $\times$ - - - $\times$ - TOLABLE
@@
7. MetaDESK 1997 (Ishii and Ullmer, 1997) Phicons $\times$ - - - $\times$ - - DATIBLE
@@
8. Build-IT 1997 (Rauterberg et al., 1997, 1998; Fjeld et al., 1999) Bricks - $\times$ - - - $\times$ - TOLABLE
@@
9. Pinwheels 1998 (Wisneski et al., 1998) Pinwheels $\times$ - - - $\times$ - - DATIBLE
10. Voodoo Dolls 1998 (Isidoro and Sclaroff, 1998) Spongy object $\times$ - - - - $\times$ - DATABLE

### v3

@@ 2.2. Blended Words
Linguistics studies the evolution of everyday language over time, including word-formation processes (e.g., derivations, abbreviations, and acronyms). Some new words are created by combining splinters (barrena2019splinters) excerpted from source words. For example, blended words concatenate a basis from a first word and a suffix from a second word, with phonemic overlap (beliaeva2016blends; connolly2013innovation; gries2004shouldnt). Linguists observed several possibilities and strategies of splinters excerption and combination that enter the definition of “blended words” that keep source words recognisable (gries2004shouldnt; gries2006cognitive). For instance, “heliport” (heli(copter) + (air)port), “motel” (mot(or) + (h)otel), and “brunch” (br(eakfast) + (l)unch) are blend words (connolly2013innovation; gries2004shouldnt). However, blends can follow some other word-formation processes. For example, clipping compounds concatenate the bases of two words (beliaeva2016blends; gries2004shouldnt; gries2006cognitive) (e.g., “fintech” (fin(ancial) tech(nology)), “sysadmin” (sys(tem) + admin(istrator)), and “hydrail” (hyd(rogen) + rail(way))). Word blending enables creating new words of only 1, 2, or 3+ syllables (connolly2013innovation), thus creating some new vocabulary that designates concepts through single words, which also facilitates communication.
Blended words already exist in Computer Science and Human–Computer Interaction: for example, “pixel” (pic(ture)s + el(ement)) and “widget” (wi(ndow) + (ga)dget) are kinds of blended words that have become common in the usual language and are referenced in standard dictionaries. In the field of tangible user interfaces, several of the aforementioned terms are blended words, such as “phicon” (ph(ysical) + icon) (ullmer1997metadesk), “phandle” (ph(ysical) + (h)andle) (ullmer1997metadesk), and “phidget” (ph(ysical) + (w)idget) (greenberg2001phidgets).
@@ 2.3. Taxonomies and Classifications
The literature has resorted to taxonomies and classifications to characterize the conceptual space of physical user interfaces. Whereas the seminal work on graspable user interfaces proposed a conceptual space with thirteen axes (fitzmaurice1995bricks), Ishii later classified tangible user interfaces through eight genres (ishii2008beyond). About two dozen frameworks organized the conceptual space along different numbers of axes mazalek2009framing, shaer2010past: Ch. 5, ullmer2022weaving: Ch. 3. For example, Fishkin’s taxonomy of physicality (or tangibility) (fishkin2004taxonomy) organized physical user interfaces as a twenty-item conceptual space along two axes: embodiment (i.e., the physical relationship between the input and the output of interaction loops) and metaphor (i.e., the digital meaning given to interaction loops with physical representations). All those frameworks were organized by types: abstracting, designing, and building; and by facets: experiences, domains, physicality, interactions, and technologies (mazalek2009framing).
@@ 3.2. Back to Tangible User Interfaces’ Concept
According to Ullmer and Ishii: “Tangible interfaces give physical form to digital information, employing physical artifacts both as representations and controls for computational media” (ullmer2000emerging). This definition’s phrasing reveals Ullmer and Ishii’s datum-centered viewpoint on tangible user interfaces (shaer2010past: p. 19). The adjective “tangible” should then be used to qualify data that becomes tangible, not to qualify artifacts or bodied objects (the latter is already “tangible” in the usual sense). This article claims we should then rather talk about “tangible data,” and no more about “tangible artifacts” or “tangible objects.” Thus, tangible data are data that become tangible through artifacts (that give physical form to data): only those “pure” artifacts must be qualified as tangible representations to match Ullmer and Ishii’s definition strictly. By extending this line of reasoning, this article proposes naming all artifacts according to their digital role—as instantiations of digital information—by providing meaning about how they represent digital entities. This way, the following section defines a what–how terminology to name physical representations and controls from a role-centered viewpoint.
@@ 3.3. A Role-Centered Terminology
The new terminology uses blended words (beliaeva2016blends; connolly2013innovation) formed from a taxonomy on “what” digital entities come into the digital world and “how”. The first part concerns what digital entities are represented or controlled by physical representations:
@@
    Operations involve triggering an event or action in the computing system (e.g., core operations (ullmer2008core) include loading or saving data, loading an app, and validation).
@@
    Finally, constraints structure the behavior of data representations, thus structuring mapped digital information. They can give meaning to artifacts by triggering actions or operations (e.g., detecting a block’s presence to trigger an action (ullmer1999mediablocks; ullmer1998mediablocks)). Broadly, “constraints” affect physical objects’ positioning (e.g., table surface) before being sensed by the computer. “Constraints” are not always computer-generated: for example, some circles where to place artifacts can be video-projected onto a surface or drawn with paint.
@@
    Those that are made “graspable” are represented by symbolic artifacts of a generic form, with a low level of embodiment. Saying tangible or graspable is not saying better or worse, but simply different gradations of embodiment (fishkin2004taxonomy) that correspond to different design and technology answers to different needs.
@@
From the seven source words mentioned above, the seven following splinters (barrena2019splinters) are excerpted: four bases from what words: “dat-,” “tol-,” “op-,” and “const-”; and three suffixes from how words: “-ible,” “-able,” and “-nible.” Several attempts were required to select those splinters by a compromise between pronunciation, syllable count, and recognition of original words in the final terms (at least for experts aware of the terminology’s construction). Blending those bases and suffixes thus provides twelve terms embedding meaning about what and how digital entities come to the physical world (see Table 1).
@@
Datum DATIBLE[”deItIb9l]‘Datum is tangible’ DATABLE[”deIt9b9l]‘Datum is graspable’ DATNIBLE[”deIt.nIb9l]‘Datum is intangible’
Tool TOLIBLE[tu:lIb9l]‘Tool is tangible’ TOLABLE[tu:l9b9l]‘Tool is graspable’ TOLNIBLE[tu:l.nIb9l]‘Tool is intangible’
Operation OPIBLE[OpIb9l]‘Operation is tangible’ OPABLE[Op9b9l]‘Operation is graspable’ OPNIBLE[Op.nIb9l]‘Operation is intangible’
Constraint CONSTIBLE[”kOnstIb9l]‘Constraint is tangible’ CONSTABLE[”kOnst9b9l]‘Constraint is graspable’ CONSTNIBLE[”kOnst.nIb9l]‘Constraint is intangible’
@@ 3.5. Four Role-Centered Tangibility Classes
  • In Class IV specimens, only operations are provided, through tangible or graspable entities (i.e., presence of opibles or opables, and data and tools are unknown or absent). This class provides specimens that are not restricted to a specific application and can thus control operations shared across applications (i.e., core operations (ullmer2008core)).
The application hallmarks’ patterns that correspond to these tangibility classes are reported in Table 2. The following section illustrates the descriptive power of the twelve terms through examples from the literature.
@@ 4.1. Specimens’ Selection
Because researchers have produced hundreds of tangible user interface specimens over the past few decades (fleck2018cladistics), the present collection is not meant to be exhaustive, but a rather representative sample. The specimen selection process is neither as rigorous as that in systematic literature reviews nor intended to provide reliable statistics. Instead, the objective is to illustrate the terminology coverage and usefulness by describing a variety of emblematic tangible user interface specimens and related physical user interfaces. Because research interests vary over the years, restricting the analysis to a specific date range could have led to the exclusion of some tangible user interface kinds; therefore, the entire research period, from the premise to the more recent work, was taken into account. The selection thus gathered various genres [^3] Seven genres of tangible user interface applications were described in 2008 (ishii2008beyond): (1) tangible telepresence, (2) tangibles with kinetic memory, (3) constructive assembly, (4) tokens and constraints, (5) interactive surfaces—tabletop TUI, (6) continuous plastic TUI, (7) augmented everyday objects, and (8) ambient media, to which actuated TUI and shape-changing TUI must now be added. of tangible user interfaces that were studied during this entire period. Finally, the selection was filtered to retain work that showed real applications. Work merely demonstrating a new technology or a new kind of tangible user interface was not retained, as its naming could vary across applications. Indeed, knowing the mapping to the digital world is necessary because being physical is not enough to be a tangible representation: the same representation can be tangible in some applications but not in others.
The selection process is summarized in Figure 1. Several sources were browsed. First, a previous state of the art (shaer2010past) and some specimens already known to the author (e.g., some specimens that were experienced during demonstration sessions at conferences over time). Second, the projects page [^4] MIT Tangible Media Group’s projects page: https://tangible.media.mit.edu/projects/, last accessed 2022-08-03. of the MIT Tangible Media Group, which is certainly the most prolific laboratory in this field (accounts for 45.5% of the final collection’s provenance). Third, to obtain even more recent work, sessions with ‘tangible’ in their title from the last ten CHI conference proceedings (from 2013 to 2022 venues). Finally, the last ten TEI conference proceedings from 2013 to 2022 were also browsed. Specimens regarding remote computer-supported conditions were rejected.
@@
State of the art (shaer2010past)
@@
Figure 1. The selection process of applicative specimens of tangible user interfaces and related physical user interfaces.
The diagram describes four successive steps: (1) Accessing sources, (2) Browsing titles, abstracts, and illustrations, (3) Extracting information from articles, and (4) Collecting specimens. The first step presents the five sources of articles. The second step sifts the articles to ensure they: (a) present a prototype, (b) answer to applicative needs, and (c) respond to single user or co-presense interaction conditions. The third step process the articles to provide applications’ terminological description, hallmarks, and genre and subgenre classification. The fourth step is iterated to pump up articles until all criteria are reached: terminological coverage, hallmarks’ orthogonality, and genres’ heterogeneity.
@@ 4.2. Corpus Description
(a) Relief: Terrain dynamic exploration (#26).
A table’s dynamic surface (datible) displays mountain reliefs colored according to elevation.
(b) SABLIER: Time navigation in cultural heritage scopes (#33).
A virtual reality cave displays the three-dimensional visualization of a cultural site. Viewpoint is set by tracking stereovision glasses. A table is placed at the cave’s center; manipulating an hourglass (opible) enables time navigation.
(c) Head Prop: Neurosurgical visualization (#5).
A rectangular plate (tolable) is placed close to a doll’s head (datible). A sizable monitor displays the resulting brain slice (datnible) from the intersection of the plate and the doll’s head.
(d) ArcheoTUI: Virtual archaeology fragments’ assembly (#23).
A monitor stands on a tabletop: two 3D fragments are displayed (datnibles); two spheres (datable) held in space control the position and orientation of the two fragments; two foot-pedals (opables) stand on the ground under the tabletop. Finally, a chair enables sitting at the tabletop.
(e) GraspDraw: 2D shape edition (#6).
Four rectangles (datnibles) are displayed on a tabletop: three small red rectangles and one sizable blue rectangle. The sizable rectangle is controlled by two bricks (tolables) placed on two opposite corners. A tray (constable) provides six functions in a row. An inkwell (constible) provides color picking between six colors: blue, green, white, yellow, red, and black. Finally, a chair enables sitting at the tabletop.
(f) GeoTUI: Cutting line selection for subsoil exploration (#24).
A map (datnible) is displayed at the center of a tabletop. A ruler (tolible) controls the cutting line (tolnible) that is displayed on the map. A button box with four buttons (opables) is placed at the right of the map. Finally, a chair enables sitting at the tabletop.
@@
1. Slot Machine 1976 (perlman1976slotmachine) and see Rows (procedure) - - - $\times$ - $\times$ - CONSTABLE
in (ullmer2022weaving: Ch. 1) Buttons (on each row) - - $\times$ - - $\times$ - OPABLE
@@
2. CAAD 3D Modelling System 1979 (aish1979caad; aish1984caad) Building Blocks $\times$ - - - - $\times$ - DATABLE
@@
3. Self-Builder Model (Segal Model) 1980 (frazer1982three; frazer1980intelligent) cited Panels $\times$ - - - $\times$ - - DATIBLE
from (sutphen2000reviving) Board - - - $\times$ - $\times$ - CONSTABLE
@@
4. Marble Answering Machine 1992 See in follmer2013inform, ishii1997tangible Machine - - - $\times$ - $\times$ - CONSTABLE
and (ullmer2022weaving: Ch. 1) Machine’s indentation (play slot) - - $\times$ - - $\times$ - OPABLE
@@
5. Head Prop 1994 (hinckley1994passiveprops) Doll’s head (brain) $\times$ - - - $\times$ - - DATIBLE
@@
6. GraspDraw 1995 (fitzmaurice1995bricks) Bricks - $\times$ - - - $\times$ - TOLABLE
@@
7. Tangible Geospace 1997 (ishii1997tangible) Phicons $\times$ - - - $\times$ - - DATIBLE
@@
8. Build-IT 1997 (rauterberg1997buildit; rauterberg1998buildit; fjeld1999camera) Bricks - $\times$ - - - $\times$ - TOLABLE
@@
9. Pinwheels 1998 (wisneski1998pinwheels) Pinwheels $\times$ - - - $\times$ - - DATIBLE
10. Voodoo Dolls 1998 (isidoro1998voodoo) Spongy object $\times$ - - - - $\times$ - DATABLE